# Fano threefolds in positive characteristic III

### v3

@@ 3.4. Conic bundles
We say that $f:X\to S$ is a conic bundle if $f:X\to S$ is a flat projective morphism of noetherian schemes such that $X_{s}$ is isomorphic to a conic on $\mathbb{P}^{2}_{\kappa(s)}$.
@@ 4.1.1. Type C
Assume that $f$ is of type $C_{2}$, i.e., every fibre of $f$ is smooth. Then $\Delta_{f}=0$ and any fibre is isomorphic to $\mathbb{P}^{1}$. Since $Y$ is a smooth projective rational surface, we have ${\operatorname{Br}}(Y)=0$ (Proposition 2.6(2)). It follows from Proposition 2.7 that $X$ is isomorphic to $\mathbb{P}(E)$ over $Y$ for some vector bundle $E$ of rank $2$. In this case, we have $\mu_{R}=2$ and $-K_{X}\cdot\ell=2$ for any fibre $\ell$. ∎
@@
Assume that $R$ is of type $C_{1}$. By Theorem 4.4, we have $\mu_{R}=1$ and $f^{-1}(y)\equiv 2\ell$ for any closed point $y$. Note that this holds even if no fibre of $f$ is smooth.
Assume that $R$ is of type $C_{2}$. Then we have $\mu_{R}=2$ and $f^{-1}(y)\equiv\ell$ for any closed point $y$. ∎
@@ 4.1.2. Type D
Let $X$ be a Fano threefold and let $R$ be an extremal ray of type $D$. Let $f\colon X\to Y$ be the contraction of $R$. Fix a closed point $y$ and $X_{y}:=f^{-1}(y)$ denotes the scheme-theoretic fibre over $y$. Then $Y\simeq\mathbb{P}^{1}$ and $X_{y}$ is a projective Goresntein surface on $X$ such that $\omega_{X_{y}}^{-1}$ is ample. Furthermore, the following hold for the length of extremal ray $\mu_{R}$ and an extremal rational curve $\ell$ of $R$.
@@ 4.1.3. Type E
Let us show (ii). By (i), $-K_{Y}$ is nef. By $-K_{Y}\cdot C>0$, $\mathcal{O}_{Y}(-K_{Y})|_{C}$ is ample. Hence $\pi^{*}(\mathcal{O}_{Y}(-K_{Y})|_{C})\simeq f^{*}(\mathcal{O}_{Y}(-K_{Y}))|_{D}$ is semi-ample. By $f^{*}(-K_{Y})=-K_{X}+D$, it follows from [CMM14, Theorem 3.2] that $f^{*}(-K_{Y})$ is semi-ample. Since $X$ and $Y$ are normal, $-K_{Y}$ is semi-ample. Thus (ii) holds. This completes the proof of Claim. ∎
@@
Thus (2) holds. By Lemma 7.2(2), we obtain
@@
Let $X$ be a primitive Fano threefold and let $R$ be an extremal ray of type $E$. Let $f\colon X\to Y$ be the contraction of $R$. Then $f$ is a birational morphism to a projective normal threefold $Y$. Furthermore, the following hold for the length of extremal ray $\mu_{R}$ and an extremal rational curve $\ell$ of $R$.
@@
The columns $\mu_{R}$ and $\ell$ can be confirmed by $(K_{X}+D)\cdot\ell=K_{D}\cdot\ell$. ∎
@@
Let $X$ be a primitive Fano threefold with an extremal ray $R$ of type $E$. Let $f:X\to Y$ be the contraction of $R$ and set $D:={\operatorname{Ex}}(f)$. Then any effective divisor $Z$ on $D$ is semi-ample.
@@
Assume that $g(D)$ is not a point, i.e., $g(D)=\mathbb{P}^{1}$. Then there exist curves $\Gamma_{1}$ and $\Gamma_{2}$ on $D$ such that $g(\Gamma_{1})$ is a point and $g(\Gamma_{2})=\mathbb{P}^{1}$. We have $g^{*}\mathcal{O}_{\mathbb{P}^{1}}(1)\cdot\Gamma_{1}=0$ and $g^{*}\mathcal{O}_{\mathbb{P}^{1}}(1)\cdot\Gamma_{1}>0$. Since the numerical equivalence classes $[\Gamma_{1}]$ and $[\Gamma_{2}]$ lie on $R\setminus\{0\}$, we can find $\lambda\in\mathbb{R}_{>0}$ satisfying $\Gamma_{1}\equiv\lambda\Gamma_{2}$, which is a contradiction. ∎
@@ 4.3. Existence of conic bundle structures
Suppose the contrary. Let $R_{1},\dots,R_{n}$ be all the extremal rays. We can find curve $\ell_{1},...,\ell_{n}$ such that
@@
Proposition 4.14.
Proof.
By Lemma 4.13, we may assume that there exists an extremal ray $R_{1}$ is of type $D$. Then we have $\rho(X)=2$ (Theorem 4.11). Let $R_{2}$ be the other extremal ray. For each $i\in\{1,2\}$, let $f_{i}:X\to Y_{i}$ be the contraction of $R_{i}$, where $Y_{1}=\mathbb{P}^{1}$. Suppose that $R_{2}$ is of type $E$ or $D$. It suffices to derive a contradiction. If $R_{2}$ is of type $D$, then Lemma 4.15(1) leads to a contradiction.
@@ 4.4. $\rho(X)\leq 3$
On the other hand, $D_{2}$ is disjoint from $f^{-1}(s)$ for a general closed point $s\in S$, and hence we get
@@
    If $f:X\to S$ is the contraction of an extremal ray of type $C$, then the following hold.
    • $\rho(X)=2$ if and only if $S\simeq\mathbb{P}^{2}$.
    • $\rho(X)=3$ if and only if $S\simeq\mathbb{P}^{1}\times\mathbb{P}^{1}$.
  2. (2)
@@
    Assume $\rho(X)=2$. Then $X$ has an extremal ray $R$ of type $C$ and we have $S\simeq\mathbb{P}^{2}$ for the contraction $f:X\to S$ of $R$.
@@
    Assume $\rho(X)=3$. Then the following hold.
    1. (i)
      Let $R$ be an extremal ray of $X$. Then $R$ is of type $C$ or $E_{1}$.
    2. (ii)
      $X$ has two extremal rays $R_{1}$ and $R_{2}$ such that $R_{1}$ is of type $C$ and $R_{2}$ is of type $C$ or $E_{1}$.
  5. (5)
    If $X$ has an extremal ray of type $C$ whose contraction $f:X\to Y$ is not generically smooth, then $p=2$, $\rho(X)=2$, and $X$ is a prime divisor on $\mathbb{P}^{2}\times\mathbb{P}^{2}$ of bidegree $(1,2)$.
@@
The assertion (1) follows from $\rho(X)=\rho(S)+1$ and Theorem 4.16. Then (1) implies (2) and (3) by Proposition 4.14. The assertion (5) follows from Theorem 4.2.
Let us show (4). Since (i) implies (ii) (Proposition 4.14), it suffices to prove (i). Suppose that there exists an extremal ray $R$ which is not of type $C$ nor $E_{1}$. Let us derive a contradiction. Since $R$ is not of type $D$, the type of $R$ is $E_{2},E_{3},E_{4}$, or $E_{5}$. Let $f:X\to Y$ be the contraction of $R$. Set $D:={\operatorname{Ex}}(f)$. There exists an extremal ray $R^{\prime}$ of type $C$ (Proposition 4.14). Let $f^{\prime}:X\to\mathbb{P}^{1}\times\mathbb{P}^{1}$ be the contraction of $R^{\prime}$. For each $i\in\{1,2\}$, consider
@@
By Corollary 4.10, we have that $g_{i}(D)$ is a point for each $i\in\{1,2\}$. Hence also $f^{\prime}(D)$ is a point. However, this is a contradiction, because any fibre of $f^{\prime}$ is one-dimensional. ∎
@@ 5. Fano threefolds with $\rho(X)=2$
In this section, we classify Fano threefolds with $\rho(X)=2$. If $X$ has an extremal ray of type C and its contraction is not generically smooth (i.e., a wild conic bunlde), then $X$ is a prime divisor on $\mathbb{P}^{2}\times\mathbb{P}^{2}$ of bidegree $(1,2)$ (Theorem 4.2). In what follows, we always assume that the contraction of type C is generically smooth.
@@ 5.1. The Picard group and the canonical divisor
    Assume that $R_{i}$ is of type $E$. Set $D_{i}:={\operatorname{Ex}}(f_{i})$. Recall that if $R_{i}$ is of type $E_{1},E_{3}$, or $E_{4}$ (resp. $E_{2}$, resp. $E_{5}$), then $r_{i}$ is defined as the largest positive integer that divides $-K_{X}+D_{i}$ (resp. $-K_{X}+2D_{i}$, resp. $-2K_{X}+D_{i}$) in ${\operatorname{Pic}}\,X$. Then $c_{2}(X)\cdot H_{i}$ satisfies the following:
@@
Let us show (1). Assume that $R_{i}$ is of type $C$. Then $f_{i}\colon X\to Y_{i}=\mathbb{P}^{2}$ is a conic bundle. Since $L_{i}$ is an ample generator of ${\operatorname{Pic}}\,\mathbb{P}^{2}$, we may assume that $L_{i}$ is a line on $Y_{i}=\mathbb{P}^{2}$. Since $H_{i}=f_{i}^{*}L_{i}$ is an effective divisor, it follows from Lemma 3.8(4) that
@@
By Corollary 3.12, it holds that
$\chi(H_{i},\mathcal{O}_{H_{i}})=1-\chi(\mathbb{P}^{2},\mathcal{O}_{\mathbb{P}^{2}}(-L_{i}))=1-\chi(\mathbb{P}^{2},\mathcal{O}_{\mathbb{P}^{2}}(-1))=1$
and
$\chi(H_{i},\mathcal{O}_{X}(H_{i})|_{H_{i}})=\chi(\mathbb{P}^{2},\mathcal{O}_{\mathbb{P}^{2}}(L_{i}))-1=\chi(\mathbb{P}^{2},\mathcal{O}_{\mathbb{P}^{2}}(1))-1=2.$
By $L_{i}^{3}=0$, we have $H_{i}^{3}=f^{*}L_{i}^{3}=0$. Since $L_{i}$ is a line on $\mathbb{P}^{2}$, it follows from Proposition 3.16 that
$(-K_{X})^{2}\cdot H_{i}=-4K_{\mathbb{P}^{2}}\cdot L_{i}-\Delta_{f_{i}}\cdot L_{i}=12-\deg(\Delta_{f_{i}}).$
To summarise, it holds that
$c_{2}(X)\cdot H_{i}$ $=6+12-0-(12-\deg(\Delta_{f_{i}}))$
$=6+\deg(\Delta_{f_{i}}).$
@@
Let us show (2). Assume that $R_{i}$ is of type $D$. Then we have $f_{i}\colon X\to Y_{i}=\mathbb{P}^{1}$. Since $L_{i}$ is an ample generator of ${\operatorname{Pic}}\,\mathbb{P}^{1}$, we may assume that $L_{i}$ is a point on $Y_{i}=\mathbb{P}^{1}$. Applying Lemma 3.8(4), we get
By Corollary 3.12, it holds that
$\chi(H_{i},\mathcal{O}_{H_{i}})=1-\chi(\mathbb{P}^{1},\mathcal{O}_{\mathbb{P}^{1}}(-L_{i}))=1-\chi(\mathcal{O}_{\mathbb{P}^{1}}(-1))=1$
and
$\chi(H_{i},\mathcal{O}_{X}(H_{i})|_{H_{i}})=\chi(\mathbb{P}^{1},\mathcal{O}_{\mathbb{P}^{1}}(L_{i}))-1=\chi(\mathcal{O}_{\mathbb{P}^{1}}(1))-1=1.$
We have $H_{i}^{3}=(f_{i}^{*}L_{i})^{3}=0$. To summarise, we obtain
@@
Let us show (3). Assume that $R_{i}$ is of type $E_{1}$. By $K_{X}=f_{i}^{*}K_{X_{i}}+D_{i}$, we obtain
$-K_{X}+D_{i}\sim-f_{i}^{*}K_{X_{i}}\sim f_{i}^{*}(r_{i}L_{i})=r_{i}H_{i}.$
@@
$c_{2}(X)\cdot H_{i}$ $=\frac{1}{r_{i}}(c_{2}(X)\cdot(-K_{X})+c_{2}(X)\cdot D_{i})$
$=\frac{1}{r_{i}}(24+c_{2}(X)\cdot D_{i}).$
Let $g(B_{i})$ be the genus of $B_{i}$. We have $d_{i}:=\deg B_{i}=(-\frac{1}{r_{i}}K_{X_{i}})\cdot B_{i}$. By [TanII, Lemma 3.21(2)], we get
@@
It holds that
$\chi(D_{i},\mathcal{O}_{D_{i}})=1-g(B_{i}),$
and
@@
$c_{2}(X)\cdot H_{i}=\frac{24}{r_{i}}+d_{i},$
@@
Assume that $R_{i}$ is of type $E_{2}$. By [TanII, Proposition 3.22], we get $\mu_{i}=2,D_{i}\simeq\mathbb{P}^{2},D_{i}^{3}=1,(-K_{X})^{2}\cdot D_{i}=4$, and $\mathcal{O}_{X}(D_{i})|_{D_{i}}\simeq\mathcal{O}_{D_{i}}(-1)$. Moreover, $\ell_{i}$ is a line on $D_{i}$. By $K_{X}\sim f_{i}^{*}K_{Y_{i}}+2D_{i}$, it holds that
$-K_{X}+2D_{i}\sim f_{i}^{*}K_{Y_{i}}\sim f_{i}^{*}(r_{i}L_{i})=r_{i}H_{i}.$
By Lemma 3.8(3), we obtain
$c_{2}(X)\cdot H_{i}$ $=\frac{1}{r_{i}}(c_{2}(X)\cdot(-K_{X})+2c_{2}(X)\cdot D_{i})$
$=\frac{1}{r}(24+2c_{2}(X)\cdot D_{i}).$
By $D_{i}\simeq\mathbb{P}^{2}$ and $\mathcal{O}_{X}(D_{i})|_{D_{i}}\simeq\mathcal{O}_{\mathbb{P}^{2}}(-1)$, it holds that $\chi(D_{i},\mathcal{O}_{D_{i}})=1$ and $\chi(D_{i},\mathcal{O}_{D_{i}}(D_{i}))=0$. It follows from Lemma 3.8(4) that

### v4

@@ 3.4. Conic bundles
We say that $f:X\to S$ is a conic bundle if $f:X\to S$ is a flat projective morphism of noetherian schemes such that $X_{s}$ is isomorphic to a conic on $\mathbb{P}^{2}_{\kappa(s)}$ for every $s\in S$.
@@ 4.1.1. Type C
Assume that $f$ is of type $C_{2}$, i.e., every fibre of $f$ is smooth. Then $\Delta_{f}=0$ and every fibre is isomorphic to $\mathbb{P}^{1}$. Since $Y$ is a smooth projective rational surface, we have ${\operatorname{Br}}(Y)=0$ (Proposition 2.7(2)). It follows from Proposition 2.8 that $X$ is isomorphic to $\mathbb{P}(E)$ over $Y$ for some vector bundle $E$ of rank $2$. In this case, we have $\mu_{R}=2$ and $-K_{X}\cdot\ell=2$ for every fibre $\ell$. ∎
@@
Assume that $R$ is of type $C_{1}$. By Theorem 4.4, we have $\mu_{R}=1$ and $f^{-1}(y)\equiv 2\ell$ for every closed point $y$. Note that this holds even if no fibre of $f$ is smooth.
Assume that $R$ is of type $C_{2}$. Then we have $\mu_{R}=2$ and $f^{-1}(y)\equiv\ell$ for every closed point $y$. ∎
@@ 4.1.2. Type D
Let $X$ be a Fano threefold and let $R$ be an extremal ray of type $D$. Let $f\colon X\to Y$ be the contraction of $R$. Fix a closed point $y$ and $X_{y}:=f^{-1}(y)$ denotes the scheme-theoretic fibre over $y$. Then $Y\simeq\mathbb{P}^{1}$ and $X_{y}$ is a projective Gorenstein surface on $X$ such that $\omega_{X_{y}}^{-1}$ is ample. Furthermore, the following hold for the length of extremal ray $\mu_{R}$ and an extremal rational curve $\ell$ of $R$.
@@ 4.1.3. Type E
Let us show (ii). By (i), $-K_{Y}$ is nef. By $-K_{Y}\cdot C>0$, $\mathcal{O}_{Y}(-K_{Y})|_{C}$ is ample. Hence $\pi^{*}(\mathcal{O}_{Y}(-K_{Y})|_{C})\simeq f^{*}(\mathcal{O}_{Y}(-K_{Y}))|_{D}$ is semi-ample. By $f^{*}(-K_{Y})=-K_{X}+D$, it follows from [CMM14, Theorem 3.2] that $f^{*}(-K_{Y})$ is semi-ample. Since $X$ and $Y$ are normal, $-K_{Y}$ is semi-ample [CT20, Lemma 2.11(4)]. Thus (ii) holds. This completes the proof of Claim. ∎
@@
Thus (2) holds. By Proposition 7.2(2), we obtain
@@
Let $X$ be a primitive Fano threefold and let $R$ be an extremal ray of type $E$. Let $f\colon X\to Y$ be the contraction of $R$. Then $f$ is a birational morphism to a projective normal threefold $Y$. Furthermore, the following hold for the length $\mu_{R}$ of the extremal ray $R$ and an extremal rational curve $\ell$ of $R$.
@@
If $R$ is not of type $E_{1}$, then the assertion follows from [TanII, Proposition 3.22]. If $R$ is of type $E_{1}$, then the assertion holds by Lemma 4.7. Note that the columns $\mu_{R}$ and $\ell$ can be confirmed by Remark 3.4.
@@
Let $X$ be a primitive Fano threefold with an extremal ray $R$ of type $E$. Let $f:X\to Y$ be the contraction of $R$ and set $D:={\operatorname{Ex}}(f)$. Then every effective divisor $Z$ on $D$ is semi-ample.
@@
Assume that $g(D)$ is not a point, i.e., $g(D)=\mathbb{P}^{1}$. Then there exist curves $\Gamma_{1}$ and $\Gamma_{2}$ on $D$ such that $g(\Gamma_{1})$ is a point and $g(\Gamma_{2})=\mathbb{P}^{1}$. We have $g^{*}\mathcal{O}_{\mathbb{P}^{1}}(1)\cdot\Gamma_{1}=0$ and $g^{*}\mathcal{O}_{\mathbb{P}^{1}}(1)\cdot{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Gamma_{2}}>0$. Since the numerical equivalence classes $[\Gamma_{1}]$ and $[\Gamma_{2}]$ lie on $R\setminus\{0\}$, we can find $\lambda\in\mathbb{R}_{>0}$ satisfying $\Gamma_{1}\equiv\lambda\Gamma_{2}$, which is a contradiction. ∎
@@ 4.3. Existence of conic bundle structures
Suppose the contrary. Let $R_{1},\dots,R_{n}$ be all the extremal rays. We can find curves $\ell_{1},...,\ell_{n}$ such that
@@
Proposition 4.15.
Proof.
By Lemma 4.13, we may assume that there exists an extremal ray $R_{1}$ of type $D$. Then we have $\rho(X)=2$ (Theorem 4.11). Let $R_{2}$ be the other extremal ray. For each $i\in\{1,2\}$, let $f_{i}:X\to Y_{i}$ be the contraction of $R_{i}$, where $Y_{1}=\mathbb{P}^{1}$. Suppose that $R_{2}$ is of type $E$ or $D$. It suffices to derive a contradiction. If $R_{2}$ is of type $D$, then Lemma 4.14(1) leads to a contradiction.
@@ 4.4. $\rho(X)\leq 3$
On the other hand, it follows from $f(D_{2})\subset E$ that $D_{2}$ is disjoint from $f^{-1}(s)$ for a general closed point $s\in S$, and hence we get
@@
    If $\rho(X)=2$, then the following hold.
    1. (2a)
      There exists an extremal ray of type $C$.
    2. (2b)
      If $f:X\to S$ is a contraction of an extremal ray of type $C$, then $S\simeq\mathbb{P}^{2}$.
@@
    If $\rho(X)=3$, then the following hold.
    1. (3a)
      There exists an extremal ray of type $C$.
    2. (3b)
      If $f:X\to S$ is a contraction of an extremal ray of type $C$, then $S\simeq\mathbb{P}^{1}\times\mathbb{P}^{1}$ and $f$ is generically smooth.
    3. (3c)
      Every extremal ray of $X$ is of type $C$ or $E_{1}$.
@@
By Proposition 4.15, (2a) and (3a) hold. Then (1), (2b), and (3b) follow from $\rho(X)=\rho(S)+1$, Theorem 4.2, and Theorem 4.16.
Let us show (3c). Suppose that $\rho(X)=3$ and there exists an extremal ray $R$ which is not of type $C$ nor $E_{1}$. Let us derive a contradiction. Note that $R$ is not of type $D$, as otherwise we would get $\rho(X)=2$ (Theorem 4.11), which contradicts $\rho(X)=3$. Hence the type of $R$ is $E_{2},E_{3},E_{4}$, or $E_{5}$. Let $f:X\to Y$ be the contraction of $R$. Set $D:={\operatorname{Ex}}(f)$. There exists an extremal ray $R^{\prime}$ of type $C$ (Proposition 4.15). Let $f^{\prime}:X\to\mathbb{P}^{1}\times\mathbb{P}^{1}$ be the contraction of $R^{\prime}$. For each $i\in\{1,2\}$, consider
@@
By Corollary 4.10, we have that $g_{i}(D)$ is a point for each $i\in\{1,2\}$. Hence also $f^{\prime}(D)$ is a point. However, this is a contradiction, because every fibre of $f^{\prime}$ is one-dimensional. ∎
@@ 5. Fano threefolds with $\rho(X)=2$
In this section, we classify Fano threefolds with $\rho(X)=2$. If $X$ has an extremal ray of type C and its contraction is not generically smooth (i.e., a wild conic bundle), then $X$ is a prime divisor on $\mathbb{P}^{2}\times\mathbb{P}^{2}$ of bidegree $(1,2)$ (Theorem 4.2). In what follows, we always assume that the contraction of type C is generically smooth.
Recall that a Fano threefold $X$ with $\rho(X)=2$ has an extremal ray of type $C$ or $E_{1}$ (Theorem 4.17). We shall treat the former (resp. latter) case in Subsections 5.2, 5.3, 5.4 (resp. Subsections 5.5, 5.6, 5.7).
@@ 5.1. The Picard group and the canonical divisor
    Assume that $R_{i}$ is of type $E$. Set $D_{i}:={\operatorname{Ex}}(f_{i})$. Then $c_{2}(X)\cdot H_{i}$ satisfies the following:
@@
Let us show (1) and (2). Assume that $R_{i}$ is of type $C$ (resp. $D$). Then we have $f_{i}\colon X\to Y_{i}=\mathbb{P}^{2}$ (resp. $f_{i}\colon X\to Y_{i}=\mathbb{P}^{1}$). Since $L_{i}$ is an ample generator of ${\operatorname{Pic}}\,\mathbb{P}^{2}$ (resp. ${\operatorname{Pic}}\,\mathbb{P}^{1}$), we may assume that $L_{i}$ is a line on $Y_{i}=\mathbb{P}^{2}$ (resp. a point on $\mathbb{P}^{1}$). Since $H_{i}=f_{i}^{*}L_{i}$ is an effective divisor, it follows from Lemma 3.8(4) that
@@
We have $H_{i}^{3}=f_{i}^{*}L_{i}^{3}=0$.
Assume that $R_{i}$ is of type $C$. It holds that
  • $\chi(H_{i},\mathcal{O}_{H_{i}})=1-\chi(\mathbb{P}^{2},\mathcal{O}_{\mathbb{P}^{2}}(-L_{i}))=1-\chi(\mathbb{P}^{2},\mathcal{O}_{\mathbb{P}^{2}}(-1))=1$ (Corollary 3.12),
  • $\chi(H_{i},\mathcal{O}_{X}(H_{i})|_{H_{i}})=\chi(\mathbb{P}^{2},\mathcal{O}_{\mathbb{P}^{2}}(L_{i}))-1=\chi(\mathbb{P}^{2},\mathcal{O}_{\mathbb{P}^{2}}(1))-1=2$ (Corollary 3.12), and
  • $(-K_{X})^{2}\cdot H_{i}=-4K_{\mathbb{P}^{2}}\cdot L_{i}-\Delta_{f_{i}}\cdot L_{i}=12-\deg(\Delta_{f_{i}})$ (Proposition 3.16).
To summarise, we get
$c_{2}(X)\cdot H_{i}=6+12-0-(12-\deg(\Delta_{f_{i}}))=6+\deg(\Delta_{f_{i}}).$
@@
Assume that $R_{i}$ is of type $D$. It holds that
  • $\chi(H_{i},\mathcal{O}_{H_{i}})=1-\chi(\mathbb{P}^{1},\mathcal{O}_{\mathbb{P}^{1}}(-L_{i}))=1-\chi(\mathcal{O}_{\mathbb{P}^{1}}(-1))=1$ (Corollary 3.12), and
  • $\chi(H_{i},\mathcal{O}_{X}(H_{i})|_{H_{i}})=\chi(\mathbb{P}^{1},\mathcal{O}_{\mathbb{P}^{1}}(L_{i}))-1=\chi(\mathcal{O}_{\mathbb{P}^{1}}(1))-1=1$ (Corollary 3.12).
To summarise, we obtain
@@
Let us show (3). Assume that $R_{i}$ is of type $E$. Set
${\small a:=\begin{cases}1\quad(\text{if $R_{i}$ is of type $E_{1},E_{3}$, or $E_{4}$})\\ 2\quad(\text{if $R_{i}$ is of type $E_{2}$})\\ \frac{1}{2}\quad(\text{if $R_{i}$ is of type $E_{5}$})\\ \end{cases}\epsilon:=\begin{cases}1\quad(\text{if $R_{i}$ is of type $E_{1},E_{2},E_{3}$, or $E_{4}$})\\ 2\quad(\text{if $R_{i}$ is of type $E_{5}$}).\end{cases}}$
Then it is easy to see that $K_{X}\equiv f_{i}^{*}K_{Y_{i}}+aD_{i}$ and
$\epsilon(-K_{X}+aD_{i})\sim-\epsilon f_{i}^{*}K_{Y_{i}}\sim r_{i}H_{i}.$
@@
$c_{2}(X)\cdot H_{i}=\frac{\epsilon}{r_{i}}(c_{2}(X)\cdot(-K_{X})+ac_{2}(X)\cdot D_{i})=\frac{\epsilon}{r_{i}}(24+ac_{2}(X)\cdot D_{i}).$
It follows from Lemma 3.8(4) that
In what follows, we apply case study to compute $c_{2}(X)\cdot D_{i}$.
Assume that $R_{i}$ is of type $E_{1}$. Let $g(B_{i})$ be the genus of $B_{i}$. We have $d_{i}:=\deg B_{i}=(-\frac{1}{r_{i}}K_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Y_{i}}})\cdot B_{i}$. By [TanII, Lemma 3.21(2)], we get
@@
It holds that $\chi(D_{i},\mathcal{O}_{D_{i}})=1-g(B_{i})$ and
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}c_{2}(X)\cdot H_{i}=\frac{\epsilon}{r_{i}}(24+ac_{2}(X)\cdot D_{i})=\frac{1}{r_{i}}(24+1\cdot r_{i}d_{i})=\frac{24}{r_{i}}+\deg B_{i},}$
@@
Assume that $R_{i}$ is of type $E_{2}$. By [TanII, Proposition 3.22], we get $\mu_{i}=2,D_{i}\simeq\mathbb{P}^{2},D_{i}^{3}=1,(-K_{X})^{2}\cdot D_{i}=4$, and $\mathcal{O}_{X}(D_{i})|_{D_{i}}\simeq\mathcal{O}_{D_{i}}(-1)$. It holds that $\chi(D_{i},\mathcal{O}_{D_{i}})=1$ and $\chi(D_{i},\mathcal{O}_{D_{i}}(D_{i}))=0$. We then get